graph the equation using the slope and y-intercept
y=5/3x+8

Answer by ewatrrr(24785)


Hi,
y=5/3x+8 is in the slope intercept form y = mx + b
slope= 5/3 & the value of y where the line crossess the y-axix is 8
Pt (0, 8)a point on the line
(-3, 3) & also points on the line according to the slope 5/3)